OpenStudy (anonymous):

Find the sum of the first 50 terms in the sequence -5, -9, -13

OpenStudy (anonymous):

Each term is represented by: -1 - 4k where k goes from 1 to 50 inclusive. \[\sum_{1}^{50} (-1 - 4k) = -50 - 4\sum_{1}^{50}k\] -50 - (4)(50)(51)/2 = -50 - (100)(51) = -50 - 5100 = -5150

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You're very welcome! @onegirl

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @onegirl